OpenStudy (anonymous):

how do i use logarithms to solve the equation 2^=14

OpenStudy (jdoe0001):

\(\bf 2^x=14\quad ?\)

OpenStudy (jdoe0001):

what do you mean by "solve"?

OpenStudy (anonymous):

i need to use logarithms to solve this equation

OpenStudy (jdoe0001):

\(\bf 2^x=14\quad ?\) <---- is that what you meant?

OpenStudy (anonymous):

yes looking to solve for what x is in logorithm form

OpenStudy (johnweldon1993):

Hint... \[\large \ln a^x => x \times \ln a\]

OpenStudy (jdoe0001):

you could also use the log cancellation rule of \(\bf log_{\color{red}{ a}}{\color{red}{ a}}^x=x\)

OpenStudy (anonymous):

thank you

OpenStudy (jdoe0001):

\(\bf log_{\color{red}{ a}}{\color{red}{ a}}^x=x \\ \quad \\ \quad \\ 2^x=14\implies log_{\color{red}{ 2}}({\color{red}{ 2}}^x(=log_{\color{red}{ 2}}14\implies x=log_214\)

OpenStudy (jdoe0001):

hmm \(\bf log_{\color{red}{ a}}{\color{red}{ a}}^x=x \\ \quad \\ \quad \\ 2^x=14\implies log_{\color{red}{ 2}}({\color{red}{ 2}}^x)=log_{\color{red}{ 2}}14\implies x=log_214\)

OpenStudy (anonymous):

this makes sense to me. thank you for your help.